OpenStudy (anonymous):

solve f(n+2) -4f(n+1) +4f(n) = n subject to f(0) = 5 and f(1) =11.

OpenStudy (jamesj):

Well, have you written out the first 10 terms or so?

OpenStudy (jamesj):

...Begin to look for patterns...

OpenStudy (anonymous):

I have to write down the f(n) =... by 2ways: generating function and use lamda method

OpenStudy (anonymous):

using generating function: I have F(x) =\[\sum_{k=0}^{\infty}f(x)x^k = f(0) +f(1)x +f(2)x^2 +f(3)x^3+......\]

OpenStudy (jamesj):

I must confess I don't know this part of mathematics. I would have to answer this question using some mix of intuition and luck. I am not familiar with how to use generating functions and the like to answer it.

OpenStudy (anonymous):

I got it. Thanks for reply my message. I have test on Friday, I will try to figure out in the book. I am near result. just need the confirmation from people who is better than me. Thanks anyway.

OpenStudy (anonymous):

I will post another question, Take a look and if you have idea please, guide me